# Overcoming Output Dimension Collapse: How Sparsity Enables Zero-shot Brain-to-Image Reconstruction at Small Data Scales

### v1

@@ 6 Quantifying hit rate and prediction error in practical sparse regression
In the previous section, we analyzed how the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ jointly determine the prediction error in sparse regression. However, the selection algorithm and the data determine these two quantities. In this section, we focus on a concrete variable selection procedure and derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Furthermore, by comparing the prediction errors by our theory with those from simulations, we confirm the accuracy of the analysis and its applicability.
Among the various variable selection methods, we focus on the simple correlation-based (akin to SIS (Fan and Lv, 2008)) as a concrete method (Figure 7). This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data. For each input vector $\mathbf{x}_{j}=\left(X_{\mathrm{tr}}\right)_{:,j}\in\mathbb{R}^{n}$ and output vector $\mathbf{y}_{k}=\left(Y_{\mathrm{tr}}\right)_{:,k}\in\mathbb{R}^{n}$ across the $n$ training samples, the model computes the sample Pearson correlation coefficient:
where $\hat{\mu}_{\mathrm{x}_{j}}=\frac{1}{n}\sum_{i=1}^{n}x_{j}^{(i)}$ and $\hat{\mu}_{\mathrm{y}_{k}}=\frac{1}{n}\sum_{i=1}^{n}y_{k}^{(i)}$ are the sample means, and $\hat{\sigma}_{\mathrm{x}_{j}}$ and $\hat{\sigma}_{\mathrm{y}_{k}}$ are the corresponding sample standard deviations. Once the model has computed the correlations $r_{jk}$ for all input-output pairs, it selects the top $p_{\mathrm{sel}}$ fraction of input variables with the highest absolute correlations $|r_{jk}|$ for each output dimension $k$. A separate ridge regression uses only the selected inputs to predict each output dimension. Applying this procedure to all $k$ yields a column-wise sparse support, which may differ across outputs.
Figure 7: Correlation-based filter. The depth labeled $n$ is the sample axis: each slice represents one training sample. For each feature in $y$, compute its correlation with all inputs in $x$ across the $n$ samples, select the top $d_{\mathrm{sel}}$ by absolute correlation, and fit the model using only the selected inputs.
Under the data generation process introduced in Section 5 (Figure 5B), we can derive a closed-form relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. We obtain the following relation for large $d$:
@@
where $\Phi$ is the cumulative distribution function of the standard normal distribution. A detailed derivation is provided in A.4. This relation defines a fixed-point equation for the hit rate $\pi$; we can estimate $\pi$ by numerically solving (14) for $\pi$ given the selection rate $p_{\mathrm{sel}}$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Figure 8 visualizes this relation for two data scales, $n/d=0.2$ and $n/d=0.8$. For a given selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$ increases as the non-zero weight ratio $a$ decreases, and a larger data scale $n/d$ leads to a larger hit rate $\pi$. In particular, when $a$ is small (i.e., the mapping is sparse), high hit rates $\pi$ are attainable even at small $p_{\mathrm{sel}}$.
Figure 8: Hit rate versus selection rate with a correlation-based filter. Each curve shows the estimated prediction error $R$ for different non-zero weight ratios $a$. The horizontal axis shows $p_{\mathrm{sel}}$, and the vertical axis shows $\pi$. Left: data scale $n/d=0.2$, right: data scale $n/d=0.8$.
@@
We next theoretically estimate the prediction error of two models: naive linear regression, which does not perform variable selection, and sparse linear regression, which performs variable selection using a correlation-based filter. Under the assumed data generation process (Figure 5B), we vary the non-zero weight ratio $a$ and the data scale $n/d$, and estimate the corresponding prediction error by using (12) and (14). We fix the noise level at $\sigma=0.1$ to facilitate comparison across conditions. For the sparse model, the selection rate $p_{\mathrm{sel}}$ is set to minimize the estimated prediction error $R$ under (12) and (14). Figure 9 summarizes the theoretical estimates. For naive linear regression, the prediction error decreases steadily with the data scale $n/d$ but remains insensitive to the sparsity $a$, so a single curve suffices; this confirms that the naive model cannot exploit the sparse structure of the data. For sparse linear regression, the prediction error decreases as the non-zero weight ratio $a$ becomes smaller. These results demonstrate that a practical sparse regression model can translate the sparse structure of the data into substantial accuracy gains and that, when the mapping is sufficiently sparse, it can achieve accurate zero-shot prediction even at small data scales.
Figure 9: The estimated prediction error from theory. The black line shows the prediction error of naive linear regression and is independent of the non-zero weight ratio $a$, so only one curve is displayed. The colored lines show the prediction error of sparse regression for different non-zero weight ratios $a$. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the estimated prediction error $R$.
To assess the accuracy and applicability of our theory, we conduct simulations under four data generation processes (Figure 10). Specifically, we consider: (A) a baseline that exactly matches the assumptions used in our analysis (Figure 5B); (B) Gaussian-distributed non-zero teacher weights that mimic uneven brain-to-feature mapping; (C) correlated signals generated with a Toeplitz covariance with parameter $\rho=0.5$ to reflect correlation structure in neural activity; and (D) input noise representing neural variability and measurement noise (see A.5 for details). Across the baseline, Gaussian weights, and input-noise settings, the theoretical estimates closely match the simulated prediction errors. For the correlated-signal setting, the theory and simulations align well within a practical range of data scales $n/d$, but we observe an increase in error at large data scales. This error increase is plausibly because a correlated signal makes variable selection more difficult, so the selection rate $p_{\mathrm{sel}}$ that is optimal in the uncorrelated case fails to select the correct variables under correlation. These results indicate that our framework provides reliable and informative guidance in practical settings.
@@
Figure 10: The simulated prediction error under four data-generating processes. Each plot shows the simulated prediction error $R$ of sparse regression under four different data-generating processes: (A) Baseline (the same as the setting in our theory), (B) Gaussian-distributed non-zero weights, (C) Correlated signals modeled by a Toeplitz covariance with $\rho=0.5$, (D) Input noise. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the simulated prediction error $R$. Each colored line shows the prediction error for a different non-zero weight ratio $a$.
@@ 7 Discussion
In this study, we theoretically analyzed how translator models behave under small data scales. We first clarified how small data scales induce output dimension collapse in naive multivariate linear regression, and we evaluate its impact on the reconstruction pipeline that employs a naive translator. We then mathematically quantify the prediction error of the sparse model by deriving formulas linking prediction error with data scale, sparsity, and other parameters. Our results show that a sparse model can achieve accurate prediction even at small data scales by exploiting the sparse brain-to-feature structure. These findings imply that sparse regression does not suffer from output dimension collapse; its gains do not arise from improving prediction within the training subspace but from enabling predictions that extend beyond that subspace.
@@
In Section 3, we derived the theoretical best prediction for naive multivariate linear regression in the translator. Although this best prediction minimizes latent-feature error, it does not always yield the most accurate reconstruction. Even so, it is evident that the output dimension collapse caused by small data scales prevents the model from accurately capturing the true latent features, thereby compromising the quality of the reconstructed images. Moreover, the best prediction analysis is valuable not only for characterizing ODC on small data scales but also for diagnosing ODC by the dataset bias described by Shirakawa et al. (2025). The calculation depends only on the training and test images (no brain data is required), so it can be performed prior to any neuroimaging experiment. Researchers can confirm and, if necessary, adjust their stimulus sets before collecting costly neural recordings.
In Section 5 and Section 6, we evaluated sparse linear regression under the idealized assumptions that model real-world data. Although real datasets do exhibit a performance boost from sparse regression, the improvement is typically more modest than in our simulations. We performed additional simulations under a range of more realistic conditions, yet some discrepancies with actual data are likely to remain. First, our simulations cover relatively high SNR regimes. Today’s brain data contains more noise (Allen et al., 2022; Hebart et al., 2023), suggesting that the effective noise level is higher than the values used in our simulations. Although our theory is applicable regardless of the noise level, obtaining a small prediction error requires sufficiently low noise. Second, the latent features used in practice may not fully satisfy the sparsity assumption. While our simulations considered several sparsity levels, it is challenging to estimate the degree of sparsity in real data. Such sparsity can be expected only when the latent features are truly “brain-like”; yet the latent features used in practice, especially fully connected layers of CNNs or the CLIP embedding, probably do not entirely meet this ideal. Improvements in neural recording denoising techniques and the design of AI feature spaces that empirically approximate the brain’s representations will likely enable zero-shot reconstruction at small data scales.
@@
Although we examined only linear translators, recent reconstruction methods sometimes deploy nonlinear architectures as the translator module (Chen et al., 2023; Scotti et al., 2023, 2024). For instance, Chen et al. (2023) employed Masked Brain Modeling, embedding structural priors of brain activity into the translator in an attempt to improve prediction accuracy. Although nonlinear networks can encode such inductive biases, output dimension collapse remains a concern. In the neural-tangent-kernel (NTK) regime (Jacot et al., 2018), nonlinear models behave like kernel linear models, so they cannot avoid output dimension collapse. Empirically, Shirakawa et al. (2025) reported that nonlinear translators used for reconstruction also experience collapse-like phenomena. A rigorous theoretical analysis of whether nonlinear translators can genuinely support zero-shot prediction remains important for future work.
@@
Furthermore, while our study focused on the translator, the generator module might partially compensate for inaccuracies in latent feature prediction. In the latest method, as described in Section 3, the improved visual appearance arises not only from replacing naive linear regression with sparse regression but also from generator-side refinements, such as stronger image priors, latent feature scaling, and carefully chosen feature inversion losses. More recently, some reconstruction pipelines began to adopt diffusion models as the generator (Ozcelik and VanRullen, 2023; Cheng et al., 2023; Scotti et al., 2024). These models enable us to generate photorealistic images even from inaccurate latent features. However, this approach carries the risk of introducing information not actually present in the brain’s representation. In other words, if the generator adds unnecessary details, the reconstructed image may stray from the accurate perceptual content encoded in neural activity. The “Recovery Check” ensures that only the information contained in the latent features is reconstructed (Shirakawa et al., 2025).
@@
Finally, our results apply to data-limited prediction problems at the interface of neuroscience and machine learning, including brain decoding and brain encoding across vision, speech, and language. For example, Brain-Score assesses brain-DNN alignment by fitting linear encoding models from model features to neural responses (Schrimpf et al., 2020). When solving regression at small data scales, output dimension collapse is a critical concern. Sparse regression is effective in avoiding collapse; however, in encoding settings, the selected DNN features tend to exhibit substantial overlap across voxels (Nonaka et al., 2021). Although we do not treat this situation here, determining how such selection concentration shapes prediction remains an important direction for future work.
7.1 Conclusion
@@
The significance of our study lies in its dual contribution: First, we show that naive linear regression suffers from output dimension collapse at small data scales, restricting generalization beyond the training data. Second, we demonstrate that sparse linear regression can achieve accurate prediction at small data scales by mathematically characterizing the prediction error. These findings provide a practical guideline for advancing image reconstruction without relying on large-scale data.
8 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT (GPT-4o, o3, 5) in order to improve readability and English translation. After using these tools/services, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
9 Acknowledgements

### v2

@@ 6 Quantifying hit rate and prediction error in practical sparse regression
We derived a prediction error expression for a broad class of sparse regression methods in terms of the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$. We next focus on a simple correlation-based filter as a concrete variable selection method and analyze how it realizes $p_{\mathrm{sel}}$ and $\pi$ under the student–teacher model. We first derive a relation between $p_{\mathrm{sel}}$, $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. We then evaluate the validity and practical range of our theoretical analysis by comparing the theoretical prediction errors with those from simulations and fMRI brain-to-image reconstruction data.
Among the various variable selection methods, we focus on a simple correlation-based filter (akin to SIS (Fan and Lv, 2008)) as a concrete procedure (Fig. 7A). This method is computationally trivial compared to wrapper or embedded approaches, making it well-suited for reconstruction pipelines that must efficiently handle large-scale data. For each output dimension, the filter selects the $d_{\mathrm{sel}}=p_{\mathrm{sel}}d$ input variables based on correlation scores (see Appendix B.5 for details). Applying this procedure separately across output dimensions yields a column-wise sparse support, which may differ across outputs.
Figure 7: Correlation-based variable selection and its hit rate under the student–teacher model. (A) Correlation-based filter for variable selection. The depth labeled $n$ is the sample axis: each slice represents one training sample. For each feature in $y$, compute its correlation with all inputs in $x$ across the $n$ samples, select the top $d_{\mathrm{sel}}=p_{\mathrm{sel}}d$ by absolute correlation, and fit the model using only the selected inputs. (B) Hit rate versus selection rate with a correlation-based filter. Each curve shows the relationship between the selection rate $p_{\mathrm{sel}}$ and the hit rate $\pi$ for different non-zero weight ratios $a$ under the student–teacher model. The horizontal axis shows $p_{\mathrm{sel}}$, and the vertical axis shows $\pi$. Left: data scale $n/d=0.2$, right: data scale $n/d=0.8$.
Under the student–teacher model (Fig. 5B), we obtain the following asymptotic relation between the selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$ as $n,d\to\infty$ with $n/d<\infty$.
@@
where $\Phi$ is the cumulative distribution function of the standard normal distribution. A detailed derivation is provided in Appendix B.6. This relation defines a fixed-point equation for the hit rate $\pi$; we can estimate $\pi$ by numerically solving Eq. 14 for $\pi$ given the selection rate $p_{\mathrm{sel}}$, the non-zero weight ratio $a$, and the data scale $\gamma=n/d$. Fig. 7B visualizes this relation for two data scales, $n/d=0.2$ and $n/d=0.8$. For a given selection rate $p_{\mathrm{sel}}$, the hit rate $\pi$ increases as the non-zero weight ratio $a$ decreases, and a larger data scale $n/d$ leads to a larger hit rate $\pi$. In particular, when $a$ is small (i.e., the mapping is sparse), high hit rates $\pi$ are attainable even at small $p_{\mathrm{sel}}$.
@@
We next theoretically estimate the prediction error of two models: naive linear regression, which does not perform variable selection, and sparse linear regression, which performs variable selection using a correlation-based filter. Under the student–teacher model (Fig. 5B), we vary the non-zero weight ratio $a$ and the data scale $n/d$, and estimate the corresponding prediction error by using Eq. 12 and Eq. 14. We fix the noise level at $\sigma=0.1$ to facilitate comparison across conditions. For the sparse model, the selection rate $p_{\mathrm{sel}}$ is set to minimize the estimated prediction error $R$ under Eq. 12 and Eq. 14. Fig. 8A summarizes the theoretical estimates. For naive linear regression, the prediction error decreases steadily with the data scale $n/d$ but remains insensitive to the sparsity $a$, so a single curve suffices; this confirms that the naive model cannot exploit the sparse structure of the data. For sparse linear regression, the prediction error decreases as the non-zero weight ratio $a$ becomes smaller. These results suggest that a sparse regression model can leverage the sparse structure of the data to achieve substantial accuracy gains and make accurate predictions even at small data scales when the mapping is sufficiently sparse.
@@
Figure 8: The theoretical and simulated prediction error. (A) The estimated prediction error in the student–teacher model. The black line shows the theoretical prediction error of naive linear regression and is independent of the non-zero weight ratio $a$, so only one curve is displayed. The colored lines show the theoretical prediction error of sparse regression for different non-zero weight ratios $a$. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the estimated prediction error $R$. (B) The simulated prediction error under four data generation processes. Each plot shows the simulated prediction error $R$ of sparse regression under four different data generation processes: (1) Baseline (the same as the setting in our theory), (2) Gaussian-distributed non-zero weights, (3) Correlated signals modeled by a Toeplitz covariance with $\rho=0.5$, (4) Input noise. The horizontal axis represents the data scale $n/d$, and the vertical axis represents the simulated prediction error $R$. Each colored line shows the prediction error for a different non-zero weight ratio $a$.
To evaluate the finite-sample accuracy of our theory and its robustness beyond the modeling assumptions, we conduct simulations under four data generation processes (Fig. 8B). Specifically, we consider: (1) a baseline that exactly matches the assumptions used in our analysis (Fig. 5B); (2) Gaussian-distributed non-zero teacher weights that mimic uneven brain-to-feature mapping; (3) correlated signals generated with a Toeplitz covariance with parameter $\rho=0.5$ to reflect the correlation structure in neural activity; and (4) input noise representing neural variability and measurement noise (see Appendix B.7 for details). Across the baseline, Gaussian weights, and input-noise settings, the theoretical estimates closely match the simulated prediction errors. For the correlated-signal setting, the theory and simulations align well within a practical range of data scales $n/d$, but we observe an increase in error at large data scales. This error increase is plausibly because a correlated signal makes variable selection more difficult, so the selection rate $p_{\mathrm{sel}}$ that is optimal in the uncorrelated case fails to select the correct variables under correlation. These results show that our theoretical error expressions remain accurate in finite samples and are robust to deviations from the modeling assumptions that reflect properties of real data. They indicate that our framework provides reliable and informative guidance in practical settings.
Figure 9: Comparison between theoretical and empirical prediction errors. (A) Theoretical prediction error of naive and sparse regression as a function of the data scale $\gamma=n/d$ under the student–teacher model. Each curve corresponds to a different non-zero weight ratio $a$, and the noise level is set to $\sigma=5$. (B) Empirical prediction errors of naive and sparse translators on the Deeprecon dataset (Shen et al., 2019) as a function of the number of training samples $n\in\{150,300,600,1200\}$. Since $d_{\mathrm{out}}$ is very large, these $n$ values correspond to small data scales $\gamma=n/d_{\mathrm{out}}<1$. Panels A and B use different error scales, so their absolute magnitudes are not directly comparable; we instead compare the relative errors of naive and sparse regression. The empirical errors show only a modest advantage of sparse regression over naive regression.
Finally, we provide an exploratory link between our theory and the empirical regime of real fMRI decoding on the Deeprecon dataset (Shen et al., 2019). We vary the number of training samples $n$ over the range $\{150,300,600,1200\}$ and evaluate the prediction error of both naive and sparse linear regression. The dimensionality of the latent features is on the order of $d_{\mathrm{out}}\approx 10^{6}$, so these training sample sizes correspond to small data scales $\gamma=n/d_{\mathrm{out}}<1$, even when accounting for redundancy in the latent representation. Existing fMRI datasets likely fall in a relatively high-noise regime. Since the trial-to-trial variability for repeated presentations is on the order of five in our noise-to-signal metric (see Appendix B.8), we use $\sigma=5$ as a convenient reference value to illustrate this regime and interpret the real-data comparison qualitatively. The true non-zero weight ratio $a$ is unknown, so we evaluate the theoretical error over a range of plausible values of $a$. Fig. 9A shows the theoretical error of naive and sparse regression, and Fig. 9B shows the empirical prediction errors on the Deeprecon dataset (see Appendix B.9 for details). Panels A and B use different error scales, so their absolute magnitudes are not directly comparable; we instead compare the relative errors of naive and sparse regression. Within the range of $\gamma$, $\sigma$, and $a$ compatible with this dataset, our analysis predicts that sparse regression yields only a modest reduction in prediction error, while a substantial part of the error remains irreducible due to noise. Consistent with this prediction, the empirical results show that sparse regression achieves a modest error reduction compared to naive regression. These results suggest that a further reduction of prediction error in existing data will require lowering measurement noise in addition to alleviating ODC.
@@ 7 Discussion
In this study, we provided a theoretical analysis of translator behavior at small data scales. We first clarified how small data scales induce output dimension collapse in naive multivariate linear regression, and we evaluated its impact on the reconstruction pipelines that employ a naive translator. We then analyzed sparse linear regression in a student–teacher model and derived an expression for the prediction error in terms of data scale, sparsity, and related parameters. Our results suggest that a sparse model can achieve accurate prediction even at small data scales by exploiting the sparse brain-to-feature structure. These findings imply that sparse regression does not suffer from ODC; its gains do not arise from improving prediction within the training subspace but from enabling predictions that extend beyond that subspace.
@@
In Section 3, we derived the theoretical best prediction for naive multivariate linear regression in the translator. This best prediction gives the lowest possible latent-feature error, so no naive translator can outperform it even under noise-free brain activity. As a caveat, the smallest latent-feature error need not correspond to the most accurate reconstructed image after applying the generator because latent-feature error and image-space error can diverge. Even so, the analysis makes clear that collapse at small data scales imposes a substantial irreducible error on the latent features, which directly limits the reconstruction quality. In addition, the best prediction analysis is valuable not only for characterizing ODC on small data scales but also for diagnosing ODC by the dataset bias described by Shirakawa et al. (2025). The calculation depends only on the training and test images (no brain data is required), so it can be performed prior to any neuroimaging experiment. Researchers can confirm and, if necessary, adjust their stimulus sets before collecting costly neural recordings.
In Section 5 and Section 6, we analyzed sparse linear regression in a student–teacher model and compared its predictions with simulations and the Deeprecon dataset. The theory predicts substantial accuracy gains when the noise level is moderate and the brain-to-feature mapping is sufficiently sparse. Current brain-to-image reconstruction data depart from this regime: the estimated noise level is high, and the sparsity of the true mapping is likely limited in practice. Under these conditions, the prediction error is dominated by noise, which accounts for the modest gains observed on real data. Moreover, strong sparsity can be expected only when the latent features are truly “brain-like”; yet the latent features used in current reconstruction studies, especially fully connected layers of CNNs or the CLIP embedding, probably do not entirely satisfy this assumption. Our results do not provide an immediate recipe for large performance improvements on existing datasets, but instead offer a quantitative guideline for future choices of measurement procedures and feature representation design: further progress will require both improved denoising of neural recordings and the construction of AI feature spaces that more closely approximate the brain’s representations, so that strong sparsity can be realized and effectively exploited.
@@
Although we examined only linear translators, several recent reconstruction methods have adopted nonlinear architectures as the translator module (Chen et al., 2023; Scotti et al., 2023; 2024). In principle, nonlinear networks can encode additional inductive biases, which may be attractive in data-scarce settings. However, simply increasing model capacity does not eliminate ODC as a concern. In the neural-tangent-kernel (NTK) regime (Jacot et al., 2018), nonlinear models behave like kernel linear models, so they cannot avoid ODC. Empirically, Shirakawa et al. (2025) reported that nonlinear translators used for reconstruction also experience collapse-like phenomena. A rigorous theoretical analysis of whether nonlinear translators can genuinely support zero-shot prediction remains important for future work.
@@
Furthermore, while our study focused on the translator, the generator module might partially compensate for inaccuracies in latent feature prediction. In the latest method of Shen et al. (2019), the improved visual appearance arises not only from replacing naive linear regression with sparse regression but also from generator-side refinements, such as stronger image priors, latent feature scaling, and carefully chosen feature inversion losses. More generally, it has been observed that a generator can recover a visually similar image even under substantial perturbations of the latent features (Onoo et al., 2025; Lee et al., 2025). This observation suggests that high image-space accuracy may be achievable even when latent-feature error remains significant. More recently, some reconstruction pipelines began to adopt diffusion models as the generator (Ozcelik and VanRullen, 2023; Cheng et al., 2023; Scotti et al., 2024). These generators can produce photorealistic images even from inaccurate latent features. However, this approach carries the risk of introducing information not actually present in the brain’s representation. In other words, if the generator adds unnecessary details, the reconstructed image may stray from the accurate perceptual content encoded in neural activity. The “Recovery Check” ensures that only the information contained in the latent features is reconstructed (Shirakawa et al., 2025).
@@
Finally, our results apply to data-limited prediction problems at the interface of neuroscience and machine learning, including brain decoding and brain encoding across vision, speech, and language. For example, Brain-Score assesses brain-DNN alignment by fitting linear encoding models from model features to neural responses (Schrimpf et al., 2020). When solving regression with small datasets, ODC is a critical concern. Sparse regression can mitigate ODC; however, in encoding settings, the selected DNN features tend to exhibit substantial overlap across voxels (Nonaka et al., 2021). Although we do not treat this situation here, understanding how such selection concentration shapes prediction remains an important direction for future work.
@@
The significance of our study lies in its dual contribution. First, we showed that naive linear regression suffers from ODC at small data scales, restricting generalization beyond the training data. Second, within a student–teacher framework, we derived a prediction error expression for sparse regression and clarified when accurate prediction is achievable at small data scales. These contributions provide a method for diagnosing ODC in existing reconstruction pipelines and offer quantitative guidelines for advancing brain-to-image reconstruction without relying on large-scale data.
Code availability.
Code is available at GitHub.
Acknowledgments